You have 8 dogs and 5 cats,what is the probability of getting 2 dogs and 2 cats.
Answer by stanbon(75887)

You have 8 dogs and 5 cats,what is the probability of getting 2 dogs and 2 cats.
----
# of ways to pick 2 dogs: 8C2 = 28
# of ways to pick 2 cats: 5C2 = 10
-----
# of pairings of 2 dogs and 2 cats: 28*10 = 280
----
# of ways to pick 4 from the 13 animals = 13C4 = 715
----
P(selecting 2dogs and 2cats) = 280/715 = 0.3916
==================================================
Cheers,
Stan H.